# WolBanking77: Wolof Banking Speech Intent Classification Dataset

### v1

@@ 4.2 Intent Detection
Intent detection is an NLP task which consists of classifying a sentence in order to detect its intention. Several contributions have been made in the field, such as Gerz et al. [2021] who published MINDS-14 which is the first training and evaluation dataset for the intent detection task using spoken data. It includes 14 intents derived from e-banking domain, with spoken examples in 14 different languages. Si et al. [2023a] presented SpokenWOZ, a large-scale speech-text dataset for spoken Task-Oriented Dialogue (TOD) in 8 domains and 249 hours of audio. Casanueva et al. [2020b] published BANKING77, a dataset made of 13,083 examples across 77 intents, in banking domain.
In this article, we want to estimate the challenging potential of our dataset to better highlight its interest for the community by evaluating downstream tasks with increasingly sophisticated ML Workflows. We first consider classic machine learning algorithms such as k-nearest neighbor (KNN), support vector machine (SVM), linear regression (LR) and naive bayes (NB) as first baselines. The results are reported in Appendix A.4. We then consider slightly more sophisticated solutions, with the LASER sentence encoder Heffernan et al. [2022] pre-trained on several languages, including Wolof, for sentence encoding, followed by standard classification models such as MultiLayer Perceptron (MLP) and Convolution Neural Networks (CNN). We report results in Appendix A.5. Then, more advanced models based on BERT are evaluated in zero-shot-classification and few-shot-classification mode, and these same models are fine-tuned on the WolBanking77 dataset for comparison. Finally, several benchmarking of recent Small Language Models (SLM) such as Llama-3.2-3B-Instruct  [^7] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct and Llama-3.2-1B-Instruct  [^8] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct are reported. To evaluate intention detection models, f1, precision and recall metrics are reported in Section 5.4.
@@ 5 Dataset evaluation & Experiments
In this section, pre-trained models on different languages, including African ones, are presented for Intent Detection and for Automatic Speech Recognition tasks. The details of the hyperparameter settings for the experiments in this section can be found in Appendix A.2. The results show the interest of the WolBanking77 dataset for the community in order to improve intent detection for Wolof and provide an approach for extension to other low-ressource languages.
@@ 5.1 Intent detection models
The next sections present and detail the pre-trained models used in the intent detection task. Experiments for zero and few shot classification are also discussed.
@@ 5.1.1 Zero-shot classification with WolBanking77
Zero-shot text classification is a natural language processing task in which a model, trained on labeled data from certain classes, is capable of classifying text from entirely new, unseen classes without additional training. In this experiments WolBanking77 provides the new, unseen datatset. To simulate Zero-shot text classification, following pre-trained models are considered:
@@
Afro-xlmr-large: Based on XLM-R-large Conneau et al. [2020] using MLM objective, this model has been published by Alabi et al. [2022] and pre-trained on 17 African languages but not Wolof, while still demonstrating in the original paper good scores for NER task on Wolof.
AfroLM_active_learning: Based on XLM-RoBERTa (XLM-R) using MLM objective Ogueji et al. [2021], this model has been published by [Dossou et al., 2022]. AfroLM_active_learning is a Self-Active Learning-based Multilingual Pretrained Language Model for 23 African Languages including Wolof.
mDeBERTa-v3-base-mnli-xnli: Based on DeBERTaV3-base He et al. [2021a, b], this model has been published by Moritz et al. [2022] and pre-trained on the CC100 multilingual dataset Conneau et al. [2020], Wenzek et al. [2020] with 100 different languages including Wolof. This model can perform Natural Language Inference (NLI) on 100 languages.
@@ 5.1.2 Few-shot classification with WolBanking77
Given the results obtained on zero-shot classification, it seems relevant to use manually annotated data in order to improve the generalization capacity of existing models on the data presented in this article, WolBanking77. All pre-trained models cited in section 5.1.1 are selected for the few-shot classification to measure the gap between large pre-trained models with and without providing few Wolof samples from WolBanking77 for fine-tuning. F1-score metrics for zero-shot text classification are reported in Section 5.4.
@@ 5.2 ASR System
State-of-the-art ASR are chosen as baselines models pre-trained on several languages with the possibility of fine-tuning them on a specific language and domain. Below is the description of the models with more details:
@@
Canary Flash: Canary Flash Żelasko et al. [2025] has been pre-trained on several languages (English, German, French, Spanish) and on various tasks such as ASR and translation. Canary Flash is based on Canary Puvvada et al. [2024] that is an encoder-decoder model whose encoder is based on the FastConformer model Rekesh et al. [2023] and the decoder on the Transformer architecture [Vaswani et al., 2017]. Canary has the particularity of concatenating tokenizers Dhawan et al. [2023] of different languages from SentencePiece. [^9] Google Sentencepiece Tokenizer In this reported work, English, Spanish, French and Wolof languages are concatenated with Canary’s Tokenizer. These tokens are then transformed into token embedding before feeding it into the Transformer Decoder. In addition to the tokens of each language, Canary uses 1152 tokens for special tokens representation. Canary has at the time of writing this article, 3 variants such as canary-1b, canary-1b-flash, and canary-180m-flash. Canary-1b-flash version is chosen for conducting experiments because of its multilingual support. The canary-1b-flash model was trained on 85K hours of speech data, including 31K public data (FLEURS Conneau et al. [2022b], CoVOST v2 Wang et al. [2020]) and the rest on private data. The Mozilla CommonVoice 12 dataset Ardila et al. [2020] is used as validation data for each language.
@@
Phi-4-multimodal-instruct: Phi-4-multimodal is a multimodal Small Language Model (SLM) supporting image, text and audio within a single model. Phi-4-multimodal has the ability to handle multiple modalities without interference thanks to the Mixture of LoRAs [Hu et al., 2021]. To support multilingual inputs and outputs, the tiktoken tokenizer  [^10] Tiktoken Tokenizer is used with a vocabulary size of approximately 200K. The model is based on a Transformer decoder Vaswani et al. [2017] and supports a context length of 128K based on LongRopE [Ding et al., 2024]. Regarding the Speech/Audio modality, several modules have been introduced including the audio encoder composed of 3 CNN layers and 24 Conformer blocks [Gulati et al., 2020]. To map the dimension (1024-d) of audio features to the text embedding space (3072-d), 2 MLP layers are used in the Audio Projector module. Note that LoRAA was used on all attention and MLP layers with a rank of 320. Phi-4-multimodal was trained on 2M hours of private speech-text pairs in 8 languages. A second post-training phase by Supervised Fine Tuning (SFT) of speech/audio data pairs was conducted. For the ASR task, it includes 20k hours of private data as well as 20k hours of public data covering 8 languages. The SFT data follows the following format:
@@
Task prompt designates the task on which the model is to be fine-tuned. Phi-4-multimodal outperforms nvidia/canary-1b-flash Żelasko et al. [2025], WhisperV3 1.5B Radford et al. [2022b], SeamlessM4T-V2 2.3B Communication et al. [2023], Qwen2-audio 8B Chu et al. [2024], Gemini-2.0-Flash Team et al. [2023] and GPT-4o OpenAI et al. [2024] on the OpenASR dataset leaderboard  [^11] https://huggingface.co/spaces/hf-audio/open_asr_leaderboard with a WER score of 6.14. Distil-whisper-large-v3.5: Distil-whisper-large-v3.5 is based on Whisper introduced by Radford et al. [2022a], which underwent training through supervised learning using 680,000 hours of labeled audio data. The authors showed that models trained with this volume of data could generalize to any dataset through zero-shot learning, meaning it can adapt without requiring fine-tuning for any specific dataset. Note that the training data encompassed 97 different languages. Several versions of Whisper followed in the last years including version 3 on which the distil-whisper-large-v3.5 model was built by knowledge-distillation following the methodology described by Gandhi et al. [2023]. Distil-whisper-large-v3.5 is trained on 98k hours of diverse filtered datasets such as Common Voice Ardila et al. [2020], LibriSpeech Panayotov et al. [2015] , VoxPopuli Wang et al. [2021] , TED-LIUM Hernandez et al. [2018], People’s Speech Galvez et al. [2021], GigaSpeech Chen et al. [2021], AMI Carletta et al. [2006], and Yodas [Li et al., 2023].
@@ 5.3 Model training
All models presented in this article are based on Pytorch library [Paszke et al., 2019]. All experiments were run on Runpod  [^12] https://www.runpod.io/, and Kaggle. [^13] https://www.kaggle.com/ P100 GPU (16GB VRAM) was used to finetune BERT base (uncased), mDeBERTa-v3-base, Afro-xlmr-large and AfroLM_active_learning. RTX 4090 GPU for AfriteVa V2. For Zero-shot and few-shot classification, RTX 2000 Ada GPU (16 GB VRAM) was used. Few-shot classification is done with SetFit huggingface library [Tunstall et al., 2022]. 2- shots refers to 2 samples per intent while 8-shots represents 8 samples per intent. All pre-trained text models are hosted on huggingface platform, they are fine-tuned using transformers library Wolf et al. [2020] except for Llama-3.2 which is fine-tuned using torchtune library [torchtune maintainers and contributors, 2024].
MPL and CNN models are trained from scratch on 200 epochs with a single 800-dimension hidden layer, ReLU activation layer and a fully connected output layer for MLP. For CNN, a Conv1d followed by a ReLU activation layer and a fully connected layer for output classification. The Wolbanking77 data was structured in the form of a prompt illustrated in Appendix A.3 in order to allow the Llama-3.2 model to generate the correct intent.
Canary Flash and Phi-4-multimodal-instruct ASR models are fine-tuned on the Wolbanking77 audio dataset using respectively the NVIDIA NEMO framework Kuchaiev et al. [2019] and huggingface Transformers on A100 SXM GPU (80GB VRAM). In the other hand, Distil-whisper-large-v3.5 is fine-tuned using the Huggingface Transformers framework on RTX 2000 Ada GPU (16 GB VRAM). All ASR models are fine-tuned on 1000 steps.
@@ 5.4 Results and Discussions
BertBaseAfroXLMRAfroLMmDeBERTa-v3AfritevaV20505010010011110.20.23311202026261717101014144646474741413535282872727979727269695555F1-scoreScores on 5k data0-shot2-shots8-shotsFT
Figure 4 shows weighted average F1-score results for zero-shot and few-shot classification. AfroXLMR model outperforms all multilingual pre-trained models on few-shot and fine-tuning (with an F1-scores of 79%) on 5k samples followed by BertBase and AfroLM. This results also show that existing models even when pre-trained on Wolof do not perform well and that our dataset provides specific challenges.
@@
BertBaseAfroXLMRAfroLMmDeBERTa-v3AfritevaV2050501001002.32.3220.40.41.51.50.50.555555757565652525050F1-scoreScores on all samples0-shotFT
Experiments have been conducted on the entire WolBanking77 dataset and reported on figure 5. Similar to previous experiments conducted on 5k sub-sample of the WolBanking77 dataset, results shows that AfroXLMR achieves slightly better scores compared to BertBase and AfroLM. These results show that intent detection for low-resource languages remains a challenging task even for state-of-the-art small language models, as shown by the results reported in table LABEL:tab:slm.
@@
A comparison is also conducted for ASR models on the audio part of WolBanking77 dataset. Special characters have been removed from the text and then reduced to lowercase. We can observe from the results with 4 hours of speech data in Table 5 that Canary-1b-flash with a wer score of 0.59% outperforms Phi-4-multimodal-instruct (wer 3.1%) and more particularly Distil-whisper-large-v3.5 (wer 4.63%). All pre-trained ASR models have been fine-tuned on the audio part of WolBanking77 dataset with 1000 steps. The results obtained following the experiments show that it is possible to obtain good results with little data, which is favorable for our WolBanking77 dataset and in low-resource language contexts.
@@ 6 Long-term Support and Future Work
Long term support includes updates to the text and audio recordings, improvements of the simulation setup script, and possible bug fixes. Free access is offered to various versions of the dataset and source code under a CC BY 4.0 license to accelerate research in the field. It is planned to add more audio recordings in various environments to make the models more robust. Text data of possible responses to each intent will be shared for possible usage in Text-To-Speech. The dataset can also be annotated for slot-filling task.

### v2

@@ 4.2 Intent Detection
Intent Detection (ID) is an NLP task that involves classifying a sentence to identify its underlying intent. Several contributions have been made in the field. For instance, Gerz et al. [2021] published MINDS-14, the first training and evaluation dataset for the ID task using spoken data. It includes 14 intents derived from e-banking domain, with spoken examples in 14 different languages. Si et al. [2023a] introduced SpokenWOZ, a large-scale speech-text dataset for spoken Task-Oriented Dialogue in 8 domains and 249 hours of audio. Casanueva et al. [2020] published BANKING77, a dataset comprising 13,083 examples across 77 intents in banking domain.
In this article, we aim to assess the challenging potential of our dataset to better highlight its relevance to the community by evaluating downstream tasks using progressively sophisticated ML Workflows. We first consider classic machine learning algorithms such as k-nearest neighbor (KNN), support vector machine (SVM), linear regression (LR) and naive bayes (NB) as initial baselines. The results are reported in Appendix 8.9. We then consider slightly more sophisticated approaches, using the LASER sentence encoder Heffernan et al. [2022] pre-trained on several languages, including Wolof, for sentence encoding, followed by standard classification models such as Multi-Layer Perception (MLP) and Convolution Neural Networks (CNN). Results are reported in Appendix 8.10. Subsequently, more advanced models based on BERT are evaluated in zero-shot-classification and few-shot-classification mode, and the same models are fine-tuned on the WolBanking77 dataset for comparison. Finally, several benchmarks of recent Small Language Models (SLM) such as Llama-3.2-3B-Instruct  [^7] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct and Llama-3.2-1B-Instruct  [^8] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct are presented. To evaluate ID models, F1, precision and recall metrics are reported in Section 5.4.
@@ 5 Dataset evaluation & Experiments
In this section, pre-trained models in multiple languages, including African languages, are presented for ID and ASR tasks. Details of the hyperparameter settings are provided in Appendix 8.6. The results demonstrate the value of the WolBanking77 dataset for the community, highlighting its potential to improve ID task in Wolof and to serve as a basis for extending approaches to other low-resource languages.
@@ 5.1 Intent detection models
The next sections present and detail the pre-trained models used in the ID task. Experiments for zero and few shot classification are also discussed.
@@ 5.1.1 Zero-shot classification with WolBanking77
Zero-shot text classification is a NLP task in which a model, trained on labeled data from specific classes, can classify text from entirely new, unseen classes without additional training. In this experiments WolBanking77 provides the new, unseen dataset. To simulate Zero-shot text classification (more details in Appendix 8.7), following pre-trained models are considered:
@@
Afro-xlmr-large: Based on XLM-R-large Conneau et al. [2020] using MLM objective, this model was published by Alabi et al. [2022] and pre-trained on 17 African languages but not Wolof, while still demonstrating in the original paper good scores for NER task on Wolof.
AfroLM_active_learning: Based on XLM-RoBERTa (XLM-R) using MLM objective Ogueji et al. [2021], this model was published by [Dossou et al., 2022]. AfroLM_active_learning is a Self-Active Learning-based Multilingual Pretrained Language Model for 23 African Languages including Wolof.
mDeBERTa-v3-base-mnli-xnli: Based on DeBERTaV3-base He et al. [2023, 2021], this model was published by Moritz et al. [2022] and pre-trained on the CC100 multilingual dataset Conneau et al. [2020], Wenzek et al. [2020] with 100 different languages including Wolof. This model can perform Natural Language Inference (NLI) on 100 languages.
@@ 5.1.2 Few-shot classification with WolBanking77
Given the results obtained in zero-shot classification, it is relevant to leverage manually annotated data to improve the generalization capacity of existing models on the dataset presented in this article, WolBanking77. All pre-trained models cited in section 5.1.1 are selected for the few-shot classification to measure the gap between large pre-trained models with and without a small number of Wolof samples from WolBanking77 used for fine-tuning. F1-score metrics for few-shot text classification are reported in Section 5.4.
@@ 5.2 ASR System
State-of-the-art ASR are selected as baseline models, pre-trained on multiple languages with the possibility of fine-tuning them for a specific language and domain. A description of these models, with additional details, is provided below:
@@
Canary Flash: Canary Flash Żelasko et al. [2025] has been pre-trained on several languages (English, German, French, Spanish) and on various tasks such as ASR and translation. Canary Flash is based on Canary Puvvada et al. [2024] that is an encoder-decoder model whose encoder is based on the FastConformer model Rekesh et al. [2023] and the decoder on the Transformer architecture [Vaswani et al., 2017]. Canary has the distinctive feature of concatenating tokenizers Dhawan et al. [2023] from different languages using SentencePiece. [^9] Google Sentencepiece Tokenizer In this work, English, Spanish, French and Wolof languages are concatenated with Canary’s Tokenizer. These tokens are then transformed into token embedding before being fed into the Transformer decoder. In addition to the tokens of each language, Canary uses 1,152 special tokens representation. At the time of writing, Canary has three variants: canary-1b, canary-1b-flash, and canary-180m-flash. Canary-1b-flash version is chosen for the experiments because of its multilingual support. The canary-1b-flash model was trained on 85K hours of speech data, including 31K hours of public data (FLEURS Conneau et al. [2023], CoVOST v2 Wang et al. [2021b]) and the remainder on private data. The Mozilla CommonVoice 12 dataset Ardila et al. [2020] was used as validation data for each language.
@@
Phi-4-multimodal-instruct: Phi-4-multimodal is a multimodal Small Language Model (SLM) supporting image, text and audio within a single model. It can handle multiple modalities without interference thanks to the Mixture of LoRAs [Hu et al., 2022]. To enable multilingual inputs and outputs, the tiktoken tokenizer  [^10] Tiktoken Tokenizer is used with a vocabulary size of approximately 200K tokens. The model is based on a Transformer decoder Vaswani et al. [2017] and supports a context length of 128K based on LongRopE [Ding et al., 2024]. For the speech/audio modality, several modules have been introduced, including an audio encoder composed of 3 CNN layers and 24 Conformer blocks [Gulati et al., 2020]. To map 1024-dimensional audio features to the 3072-dimensional text embedding space, 2 MLP layers are used in the Audio Projector module. Note that LoRAA was used to all attention and MLP layers with a rank of 320. Phi-4-multimodal was trained on 2M hours of private speech-text pairs in 8 languages. A second post-training phase using Supervised Fine Tuning (SFT) on speech/audio data pairs was then conducted. For the ASR task, this included 20k hours of private data and 20k hours of public data across 8 languages. The SFT data follows the format below:
<|user|><audio>{task prompt}<|end|><|assistant|>{label}<|end|>
The task prompt designates the specific task on which the model is to be fine-tuned.
Phi-4-multimodal outperforms nvidia/canary-1b-flash Żelasko et al. [2025], WhisperV3 1.5B Radford et al. [2023], SeamlessM4T-V2 2.3B Communication et al. [2023], Qwen2-audio 8B Chu et al. [2024], Gemini-2.0-Flash Team et al. [2023] and GPT-4o OpenAI et al. [2024] on the OpenASR dataset leaderboard  [^11] https://huggingface.co/spaces/hf-audio/open_asr_leaderboard with a WER score of 6.14.
@@
Distil-whisper-large-v3.5: Distil-whisper-large-v3.5 is based on Whisper introduced by Radford et al. [2023], which was trained through supervised learning on 680,000 hours of labeled audio data. The authors demonstrated that models trained with this scale could generalize to any dataset through zero-shot learning, meaning they can adapt without requiring fine-tuning for specific datasets. The training data covered 97 different languages. Several versions of Whisper have been released in recent years, including version 3 on which the distil-whisper-large-v3.5 model was built by knowledge-distillation following the methodology described by Gandhi et al. [2023]. Distil-whisper-large-v3.5 was trained on 98k hours of diverse filtered datasets such as Common Voice Ardila et al. [2020], LibriSpeech Panayotov et al. [2015] , VoxPopuli Wang et al. [2021a] , TED-LIUM Hernandez et al. [2018], People’s Speech Galvez et al. [2021], GigaSpeech Chen et al. [2021], AMI Carletta et al. [2006], and Yodas [Li et al., 2023].
@@ 5.3 Model training
All models presented in this article are based on Pytorch library [Paszke et al., 2019]. All experiments were run on Runpod  [^12] https://www.runpod.io/, and Kaggle. [^13] https://www.kaggle.com/ P100 GPU (16GB VRAM) was used to finetune BERT-Base, mDeBERTa-v3-base, Afro-xlmr-large and AfroLM_active_learning. RTX 4090 GPU for AfriteVa V2. For Zero-shot and few-shot classification, RTX 2000 Ada GPU (16 GB VRAM) was used. Few-shot classification is done with SetFit huggingface library [Tunstall et al., 2022]. 2-shots refers to 2 samples per intent while 8-shots represents 8 samples per intent. All pre-trained text models are hosted on huggingface platform, they are fine-tuned using transformers library Wolf et al. [2020] except for Llama-3.2 which is fine-tuned using torchtune library [torchtune maintainers, 2024].
MPL and CNN models were trained from scratch for 200 epochs. The MLP model consists of a single 800-dimensional hidden layer, a ReLU activation layer, and a fully connected output layer. For CNN, a Conv1d layer followed by a ReLU activation layer and a fully connected layer for output classification. The Wolbanking77 data was structured in the form of a prompt, as illustrated in Appendix 8.8, to enable the Llama-3.2 model to generate the correct intent.
Canary Flash and Phi-4-multimodal-instruct ASR models are fine-tuned on the Wolbanking77 audio dataset using the NVIDIA NEMO framework Kuchaiev et al. [2019] and Huggingface Transformers library, respectively, on an A100 SXM GPU (80GB VRAM). In contrast, Distil-whisper-large-v3.5 is fine-tuned using the Huggingface Transformers framework on an RTX 2000 Ada GPU (16 GB VRAM). All ASR models are fine-tuned for 1,000 steps.
@@ 5.4 Results and Discussions
BERT-BaseAfroXLMRAfroLMmDeBERTa-v3AfritevaV20505010010011110.20.23311202026261717101014144646474741413535282872727979727269695555F1-scoreScores on 5k data0-shot2-shots8-shotsFT
Figure 4 shows the weighted average F1-score results for zero-shot and few-shot classification. AfroXLMR model outperforms all multilingual pre-trained models in few-shot and fine-tuning settings (with an F1-scores of 79% on 5k samples), followed by BERT-Base and AfroLM. This results also show that existing models, even when pre-trained on Wolof, do not perform well and that our dataset presents specific challenges.
@@
BERT-BaseAfroXLMRAfroLMmDeBERTa-v3AfritevaV2050501001002.32.3220.40.41.51.50.50.555555757565652525050F1-scoreScores on all samples0-shotFT
Experiments were conducted on the entire WolBanking77 dataset and reported in figure 5. Similar to the experiments performed on the 5k sub-sample of WolBanking77, results shows that AfroXLMR achieves slightly better scores compared to BERT-Base and AfroLM. These findings highlight that ID task for low-resource languages remains a challenging task, even for state-of-the-art small language models, as shown in Table LABEL:tab:slm.
@@
A comparison was also conducted for ASR models on the audio part of WolBanking77 dataset. Special characters were removed from the text, and all text was converted to lowercase. We can observe from the results with 4 hours of speech data in Table 5 that, Canary-1b-flash with a WER score of 0.59% outperforms Phi-4-multimodal-instruct (WER 3.1%) and more particularly Distil-whisper-large-v3.5 (WER 4.63%). All pre-trained ASR models were fine-tuned on WolBanking77 audio dataset for 1000 steps. The results indicate that strong performance can be achieved with relatively little data, which is promising for WolBanking77 and other low-resource language contexts.
@@ 6 Long-term Support and Future Work
Long term support includes updates to the text and audio recordings, improvements to the simulation setup script, and potential bug fixes. Free access is provided to various versions of the dataset and source code under a CC BY 4.0 license to accelerate research in the field. It is planned to add more audio recordings in diverse environments to enhance model robustness. Text data corresponding to possible responses for each intent will be shared for potential use in Text-To-Speech applications. Additionally, the dataset can be annotated for slot-filling tasks.